# Privacy-Aware Detection of Fake Identity Documents: Methodology, Benchmark, and Improved Detection Methods (FakeIDet2)

### v1

@@ 1 Introduction
The veracity of digital information is one of the great challenges of society today [2, 37]. With the great advances made in the field of GenerativeAI [24], it is possible to synthesize non-existent content [20], or to modify existing one [28], using simple and fast toolbox available on the internet. These methods can be used for good purposes, e.g., correct biases [29] or improve performance in some scenarios where data scarcity is present [21, 31], but they have their downsides, as they can also be used for malicious purposes such as DeepFakes [34, 35] or misinformation [18]. Evidences of the latter have been found in recent news, where tampered [^2] https://www.nytimes.com/2025/02/13/nyregion/students-high-tech-fake-ids.html or synthetic [^3] https://www.404media.co/inside-the-underground-site-where-ai-neural-networks-churns-out-fake-ids-onlyfake/ documents were used for underage alcohol purchase or bypass remote verification systems such as Know Your Customer (KYC) to access digital services such as crypto-exchanges or digital banking.
@@
In order to advance in this challenging topic, the present study focuses on Presentation Attack Detection (PAD) [15, 11] on Identity Documents (IDs). Concretely, as can be seen in Fig. 1, we focus on the most popular physical attacks in real-world scenarios: i) print attacks are physical fake IDs created using an image or a scanned Identity Document (ID), which is printed on glossy paper to resemble the texture and appearance of real IDs; ii) screen attacks, which consist in taking a picture of an ID displayed on a digital screen such as a laptop, tablet, or smartphone; and iii) composite attacks, which do not require to tamper an entire document, but only small portions, such as date of birth, family name or the portrait image from the ID owner. In the particular case of physical composite attacks, the impostor first prints fields with fake information on colored paper that resembles the colors of the target ID, and crop each field. After that, these crops are lied on the ID document, occluding the original information with the fake one. All these kinds of attacks, a.k.a Presentation Attack Instruments (PAIs), are currently being used by attackers to surpass KYC procedures, which explicitly require capturing live images of the user’s ID using an smartphone camera, granting them access to digital services using fake identities.
@@
Figure 1: Examples of a real ID and popular physical attacks (print, screen and physical composite) available in our proposed FakeIDet2-db. *Note ‐ The personal information is redacted in this figure to protect the privacy of the subject.
@@
Several studies have preliminary analyzed the problem of fake ID detection, proposing very valuable ideas and resources [5, 8, 13]. However, there are several limitations in the field that must still be covered to properly advance in this research direction. First, there are no public databases with official (a.k.a. bona fide) IDs acquired in variable conditions, nor a standard benchmark that evaluates the performance of fake ID detection systems on real-world scenarios. So far, all public databases in the literature consider as “real” IDs those created by researchers in laboratory conditions, not official governments. From now on, we refer to these non-official IDs created by research groups as “real”. As a result, these “real” IDs lack important information and patterns such as watermarks, not representing therefore a fidelity scenario of the problem. Some efforts have been recently carried out in this direction through the organization of international challenges such as DeepID [^4] https://deepid-iccv.github.io/ (although the evaluation was private and is not publicly available yet) highlighting a considerable gap in performance between using “real” IDs and official ones (i.e., 0.99 vs. 0.71 F1-score for the winner of the challenge). The main reason for considering this “real” scenario is due to privacy concerns, as some fields of the ID contain sensitive information. This is one of the key motivations for the present study, to research possible solutions to previous concerns through the exploration of realistic scenarios in which research institutions can collaborate with, for example, governments or banks to advance in this challenging research direction.
@@
Figure 2: Our proposed framework for privacy-aware fake ID detection. Current scenario (top) carries privacy concerns regarding data regulations, and need for complete ID share between ID Holders and AI Researchers. Our framework proposes an scenario where sequences of non-ordered patches extracted from different levels of anonymized real IDs can be used to preserve ID owners’ sensitive information for data sharing and inference in both in-house or outsourced schemes.
@@
In the present study, we hypothesize that using patches from IDs, instead of the whole ID, to train fake ID detectors might provide benefits from both performance and privacy points of view in real-world scenarios. Fig. 2 provides a graphical representation of the current scenario considered nowadays for the problem of fake ID detection (top), versus our proposed privacy-aware proposed scenario (bottom). In both scenarios there are two main actors: ID Holders (e.g., government, police, bank, etc.), which provide digital services to citizens and own large-scale datasets of real IDs, and the AI Researchers (e.g., research institutions), which have the experience and technology to develop fake ID detectors, but not the data. The current scenario based on sharing the whole ID raises privacy concerns, limiting the development of robust fake ID detectors. Also, there are important concerns in terms of attacks, e.g., if the attackers are able to access to the model trained on whole ID private data, they could use techniques that retrieve the training data with small prior information about the data distribution [12, 40].
@@
Our proposed privacy-aware scenario is based on the premise that an individual patch contains much less sensitive information than the whole ID. In addition, in order to provide flexibility to the particular restrictions of each ID Holder, we explore different anonymization levels, removing patches from sensitive regions as desired. For example, ID Holders may request AI Researchers to train a fake ID detector but they do not want to share any personal information about the ID owners. As a result, ID Holders can anonymize the sensitive sections from an ID, depending on their preferences for the task at hand, and extract patches from the areas which hold non-anonymized information. Additionally, to increase privacy, the extracted patches can be randomized regarding the order in which they are shared with the AI Researchers, further preventing reconstruction. This flexibility allows a fruitful collaboration between ID Holders and AI Researchers in order to train fake ID detectors under any ID Holder restrictions.
@@
Once the fake ID detector is trained by the AI Researchers, the ID Holders can decide whether inference is carried out inside the AI Researchers infrastructures or internally, as seen in Fig. 2. For example, if the ID Holders want to perform in-house inference using sensitive information that they only have access to, the AI Researchers can transfer the ID Holders the fake ID detector previously trained using anonymized IDs. However, if the ID Holders do not have such resources, they can rely on the AI Researchers for inference, since the ID Holders always control the amount of sensitive information shared for each ID. The proposed scenario can also be applicable to the research community in order to advance in this challenging topic, since research groups can share patches of IDs, including different levels of anonymization as desired.
@@
  • We explore novel approaches in order to facilitate the collaboration between ID Holders and AI Researchers in real-world fake ID detection scenarios. Concretely, instead of feeding the fake ID detectors with the whole ID, as commonly done in the literature, we propose to explore privacy-aware scenarios based on different patch sizes ($\times64$ and $\times128$) and anonymization configurations (non-, pseudo-, and fully-anonymized ID). This proposal increases flexibility to both ID Holders and AI Researchers for both training and inference stages, as shown in Fig. 2. Depending on the particular restrictions of the ID Holder, different setups can be selected to optimize both privacy and performance.
@@
  • We provide a novel database, FakeIDet2-db, which contains official real data, along with print, screen, and physical composite attacks. To the best of our knowledge, FakeIDet2-db is the first database containing physical composite forgeries, as previous ones considered digital composites. In total, FakeIDet2-db comprises 922,057 patches extracted from 2,000 ID images (pseudo-, and fully-anonymized configurations) from 47 different Spanish IDs. In addition, in order to increase the variability during the acquisition, we consider three smartphones for different end-users, including different hardware and software: iPhone 15 (high-end), Xiaomi Mi 9T Pro (medium-end) and Redmi 9C NFC (low-end). To add further variability, images are taken at 3 different heights (10 cms, 12.5 cms and 15 cms) with respect to the ID document and also different illumination conditions: no-light with the camera flash activated, dim-light with/without the camera flash activated and bright-light with/without the camera flash activated.
@@
  • We propose a novel privacy-aware fake detector, FakeIDet2, which introduces two novel learnable modules: the Patch Embedding Extractor and the Patch Embedding Fusion. The Patch Embedding Extractor leverages DINOv2 foundation model to learn representations of individual patches from real/fake IDs, using an optimized AdaFace loss function [17] through adaptable class-weights, which evolve as the training progresses. Once the embeddings are obtained, they are forwarded jointly into the Patch Embedding Fusion, based on Multi-Head Self-Attention (MHSA) [36], learning relations between patch representations to classify each ID.
@@
  • We have developed a standard, reproducible benchmark on which we evaluate the performance of our proposed FakeIDet2 in 3 different ID databases (DLC-2021 [30], KID34K[27], and Benalcazar et al. synthetic database from [5]), considering both physical and synthetic attacks, with the goal of assessing the performance in realistic out-of-distribution data. The proposed FakeIDet2-db and benchmark will be publicly available for reproducibility reasons in our GitHub repository [^5] https://github.com/BiDAlab/FakeIDet2-db, allowing a straightforward comparison of recent approaches with the state of the art.
@@
A preliminary analysis of this privacy-aware scenario was conducted in [23]. The present article significantly improves [23] in the following aspects: i) we present our novel publicly available FakeIDet2-db, which increases the number of unique real IDs (47 subjects) along with more acquisition devices (3), illumination (3) and height (3) conditions, resulting in 922,057 patches extracted from 2,000 ID images (pseudo-, and fully-anonymized configurations); ii) we include physical composite attacks in our novel FakeIDet2-db. To the best of our knowledge, this is the first database that includes this kind of Presentation Attack Instrument (PAI); iii) we propose a novel approach for fake ID detection, FakeIDet2, that overcomes the limitations regarding patch-level fusion encountered in FakeIDet [23]. While FakeIDet performs a simple mean of scores assigned to each individual patch for IDs level prediction, FakeIDet2 comprises two learnable modules (i.e., Patch Embedding Extractor and Patch Embedding Fusion) to leverage the fusion of learned patch representations, depending on the visual characteristics from real/fake IDs. The representations at patch level are obtained through an optimized version of AdaFace [17], including dynamic class weights, which accounts for the severe class imbalance present in the databases for the different PAIs. The fusion of the representations is performed via MHSA, which ponders the embeddings values based on the relations between them, to then compress them into a single embedding, via attention pooling based on [4]. The impact of these improvements can be seen in the results (Sec. 6); and iv) in order to advance in this research field, we propose a standard public benchmark considering a wide variety of physical and synthetic attacks from different databases in the literature, allowing a direct comparison of recent approaches with the state of the art.
@@
The rest of the paper is organized as follows. Sec. 2 provides an overview of the fake ID detection problem. In Sec. 3 we present our FakeIDet2-db, delving into the acquisition process for capturing real IDs and their fake counterparts. Sec. 4 introduces our novel privacy-aware fake detection method, FakeIDet2. Sec. 5 proposes a reproducible experimental protocol and standard benchmark, considering experiments in both intra- and cross-database scenarios. Finally, Sec. 6 analyzes the performance of our proposed fake detector using the proposed experimental protocol and standard benchmark and Sec. 7 draws the final conclusions.
@@ 2 Related Work
This section provides an overview of the fake ID detection problem. Sec. 2.1 provides a revision of databases considered in the field. Sec. 2.2 focuses on representative fake ID detectors and recent international challenges. Finally, we summarize key aspects in Sec. 2.3.
@@
Table 1: Summary of private and public databases considered in the literature. For completeness, we also indicate whether the databases used to train fake ID detectors are available or not for public research.
@@ 2.1 Databases
Table 1 provides a summary of the most relevant public/private databases in the field of fake ID detection. One of the first family of databases was the MIDV-family [1, 8, 7]. With the original purpose of Optical Character Recognition (OCR), Bulatov et al. synthetically created a set of physical fake ID, passports and driving licenses. They used several digital templates from multiple countries which they filled with names and addresses from Wikipedia and artificially generated faces. As different versions from the database were released, the number of fake samples were increased. DLC-2021 database [30] used the documents of the MIDV-family to create PAI s: color print, gray print and screen. The authors also provided “real” samples, although they should be considered fake samples as they are just high-quality prints built by them, not official ID. A similar work is presented in the KID34K dataset [27], where Park et al. introduced a dataset of 34,000 images from 82 Korean IDs and driving licenses. As in the MIDV-family, the authors claimed to release “real” samples, but they were also built by the authors, not official documents. In order to prove the quality of the “real” samples, they trained a Convolutional Neural Network (CNN) using the KID34K dataset, forwarding the whole ID. Evaluation was performed using both official Korean IDs and “real” documents. After that, the authors applied a dimensionality reduction technique to visualize the model’s embeddings in 2D, where they showed that embeddings from both official and “real” generated IDs from the KID34K database were similarly distributed in the embedding space. These results raise questions regarding security in the fake ID problem, since fake IDs may be generated with such resemblance to real ones, that deep-learning models may not be able to extract distinguishable features for proper PAD. Soares et al. [32] introduced the Brazilian Identity Document (BID) database with a total of 28,800 images from Brazilian IDs that have been digitally altered via different algorithms that remove, blur and replace different personal information from the real IDs, to comply with data privacy regulations. In [13], Gonzalez and Tapia proposed a more sophisticated attack procedure, i.e., printing the digital copy over a Poli-Vinyl Chloride (PVC) card, which resembles even more the appearance of real IDs.
@@
Regarding synthetic data, Benalcazar et al. proposed in [5] to use Generative Adversarial Networks (GANs) to produce synthetic Chilean IDs. The proposed GAN was trained using only real Chilean ID samples. Although the data synthesized by the authors is not strictly real, this was presented as a good idea to cover partially the lack of real data, e.g., as a data augmentation strategy in training. Another synthetic database is IDNet [39], where Xie et al. proposed a database that comprises synthetic copies of IDs from 10 countries and 10 states from the United States of America. Given an ID template, portrait images and data from the owners were synthetically generated, with special care that the fields matched with the gender and age of the synthetic portrait image. Additionally, the authors created six different types of digital manipulations, including inpainting, field cropping or face morphing. A total of 837,060 IDs were generated, making it the biggest and most diverse to this day, although no real data is included. Markham et al. [19] developed Open-Set, a database that leverages neural style transfer, via GANs-based models to transfer textures from print and screen attacks to “real” images from the MIDV-2020 and DLC-2021 databases, creating even more samples of attacks.

### v2

# Privacy-Aware Detection of Fake Identity Documents: Methodology, Benchmark, and Improved Algorithms (FakeIDet2)

@@ 1 Introduction
Analyzing the veracity of digital information is one of the great challenges facing society today [2, 37]. With the great advances made in the field of Generative AI, it is possible to synthesize non-existent content or to modify existing content [21, 26, 36], using simple and fast tools easily available on the Internet. These methods can be used for good purposes, for example, correct biases [29] or improve performance in some scenarios where data scarcity is present [22, 32], but they have their downsides, as they can also be used for malicious purposes, such as creating DeepFakes that are harmful [35, 36] or misinformation [18]. Evidence of the latter has been found in recent news, where tampered documents [^2] https://www.nytimes.com/2025/02/13/nyregion/students-high-tech-fake-ids.html or synthetic [^3] https://www.404media.co/inside-the-underground-site-where-ai-neural-networks-churns-out-fake-ids-onlyfake/ were used for the purchase of underage alcohol or to bypass remote verification systems such as Know Your Customer (KYC) to access digital services such as crypto exchanges or digital banking.
@@
In order to advance in this challenging topic, the present study focuses on Presentation Attack Detection (PAD) [12, 16] on Identity Documents (IDs)  [^4] Other related works define ID as “Identity”. Here we define ID as “Identity Document”.. Concretely, as can be seen in Fig. 1, we focus on the most popular physical attacks in real-world scenarios: i) print attacks are physical fake IDs created using an image or a scanned Identity Document (ID), which is printed on glossy paper to resemble the texture and appearance of real IDs; ii) screen attacks, which consist in taking a picture of an ID displayed on a digital screen such as a laptop, tablet, or smartphone; and iii) composite attacks, which do not require to tamper an entire document, but only small portions, such as date of birth, family name or the portrait image from the ID owner. In the particular case of physical composite attacks, the impostor first prints fields with fake information on colored paper that resembles the colors of the target ID, and then crops each field. After that, these crops are lied on the ID document, occluding the original information with the fake one. All these kinds of attacks, a.k.a. Presentation Attack Instruments (PAIs), are currently being used by attackers to surpass KYC procedures, which explicitly require capturing live images of the user ID using a smartphone camera, granting them access to digital services using fake identities.
@@
Figure 1: Examples of a real ID and popular physical attacks (print, screen, and physical composite) available in our proposed FakeIDet2-db. The personal information is redacted in this figure to protect the privacy of the subject.
@@
Several studies have preliminarily analyzed the problem of fake ID detection, proposing very valuable ideas and resources [5, 7, 15]. However, there are several limitations in the field that need to be covered in a proper way to advance this research direction. First, there are no public databases with official (a.k.a. bona fide) IDs acquired under variable conditions, nor a standard benchmark that evaluates the performance of fake ID detection systems in real-world scenarios. So far, all public databases in the literature consider “real” IDs some created by researchers under laboratory conditions, not official governments. From now on, we refer to these non-official IDs created by research groups as “real” (with quotation marks). As a result, these “real” IDs lack important information and patterns such as watermarks, not representing therefore a fidelity scenario of the problem. Some efforts have recently been made in this direction through the organization of international challenges such as DeepID [^5] https://deepid-iccv.github.io/ (although the evaluation was private and is not publicly available yet) highlighting a considerable gap in performance between using “real” IDs and the official ones (0.99 vs 0.71 F1 score for the winner of the challenge). The main reason for considering this “real” scenario is due to privacy concerns, as some fields of the IDs contain very sensitive information. These observations from state-of-the-art research trigger our key motivation for the present study: to research privacy-aware methods where more realistic data than the commonly used “real” data can be used to develop better fake ID detectors based on machine learning, while keeping the data handling as private as possible.
@@
Figure 2: Our proposed framework for privacy-aware fake ID detection. The current scenario (top) raises privacy concerns regarding data regulations, and the need for complete ID sharing between ID Holders and AI Researchers. Our framework proposes a scenario where collections of non-ordered patches extracted from different levels of anonymized IDs can be used to preserve the sensitive information of ID owners during data sharing for training and inference in both in-house and outsourced schemes.
@@
In the present study, we hypothesize that using patches from IDs, instead of the whole ID, to train fake ID detectors might provide benefits from both the performance and privacy points of view in real-world scenarios. Fig. 2 provides a graphical representation of the current scenario considered today for the problem of fake ID detection (top), versus our proposed privacy-aware scenario (bottom). In both scenarios, there are two main actors: ID Holders (e.g., government, police, bank, etc.), which provide digital services to citizens and own large-scale datasets of real IDs, and the AI Researchers (e.g., research institutions), which have the experience and technology to develop fake ID detectors, but not the data. The current scenario based on sharing the whole ID raises privacy concerns, limiting the development of robust fake ID detectors. Also, there are important concerns in terms of attacks, e.g., if the attackers are able to access the model trained on whole ID private data, they could use techniques that retrieve training data with small prior information about the data distribution [9, 13, 40].
@@
Our proposed privacy-aware scenario is based on the premise that an individual patch contains much less sensitive information than the whole ID. In addition, to provide flexibility to the particular restrictions of each ID Holder, we explore different anonymization levels, removing patches from sensitive regions as desired. For example, ID Holders may request AI Researchers to train a fake ID detector but they may not want to share any personal information about the ID owners. As a result, ID Holders can anonymize the sensitive sections from an ID, depending on their preferences for the task at hand, and extract patches from the areas that hold non-anonymized information. Additionally, to increase privacy, the extracted patches can be randomized regarding the order in which they are shared with the AI Researchers, further preventing reconstruction. This flexibility allows for fruitful collaboration between ID Holders and AI Researchers to train fake ID detectors in any ID Holder restrictions.
@@
Once the fake ID detector is trained by the AI Researchers, the ID Holders can decide whether inference is carried out inside the AI Researchers infrastructures or internally, as seen in Fig. 2. For example, if ID Holders want to perform in-house inference using sensitive information that they only have access to, AI Researchers can transfer the fake ID detector previously trained using anonymized IDs to ID Holders. However, if ID Holders do not have such resources; they can rely on the AI Researchers for inference, since the ID Holders always control the amount of sensitive information shared for each ID. The proposed scenario can also be applicable to the research community in order to advance in this challenging topic, since research groups can share patches of IDs, including different levels of anonymization as desired.
@@
  • We propose and explore a new methodology to facilitate the collaboration between ID Holders and AI Researchers for advancing real-world fake ID detection. Concretely, instead of feeding the fake ID detectors with the whole ID, as is commonly done in the literature, we explore privacy-aware scenarios based on different patch sizes ($\times64$ and $\times128$) and anonymization configurations (non-, pseudo-, and fully-anonymized ID). This proposal increases flexibility to both ID Holders and AI Researchers for both the training and inference stages, as shown in Fig. 2. Depending on the particular restrictions of the ID Holder, different setups can be selected to trade-off and optimize both privacy and performance.
@@
  • We provide a new database, FakeIDet2-db, which contains official real data, along with print, screen, and physical composite attacks. To our knowledge, FakeIDet2-db is the first database containing physical composite forgeries, as previous ones considered digital composites. In total, FakeIDet2-db comprises 922,057 patches extracted from 2,000 ID images (pseudo- and fully-anonymized configurations) from 47 different Spanish IDs. In addition, to increase the variability during the acquisition, we consider three smartphones for different end-users, including different hardware and software: iPhone 15 (high-end), Xiaomi Mi 9T Pro (medium-end) and Redmi 9C NFC (low-end). To add further variability, images are taken at 3 different heights (10 cms, 12.5 cms, and 15 cms) with respect to the ID and also different illumination conditions: no-light with the camera flash activated, dim-light with/without the camera flash activated, and bright-light with/without the camera flash activated.
@@
  • We propose a novel privacy-aware fake ID detector, FakeIDet2, which introduces two new learnable modules: the Patch Embedding Extractor and the Patch Embedding Fusion. The Patch Embedding Extractor leverages the DINOv2 foundation model to learn representations of individual patches from real/fake IDs, using an optimized AdaFace loss function [17] through adaptable class weights, which evolve as training progresses [14]. Once the embeddings have been obtained, they are merged jointly in the Patch Embedding Fusion module.
@@
  • We have developed a standard reproducible benchmark on which we evaluate the performance of our proposed FakeIDet2 in 3 different ID databases (DLC-2021 [30], KID34K [28], and Benalcazar et al. synthetic database of [5]), considering both physical and synthetic attacks, with the goal of assessing the performance in realistic out-of-distribution data. The proposed FakeIDet2-db and benchmark will be publicly available for reproducibility reasons in our GitHub repository [^6] https://github.com/BiDAlab/FakeIDet2-db, allowing a straightforward comparison of recent approaches with the state of the art.
@@
A preliminary analysis of this privacy-aware scenario was conducted in [25]. The present article significantly improves [25] in the following aspects: i) We present our new publicly available FakeIDet2-db, which increases the number of unique real IDs (47 subjects) along with triple number of acquisition devices (now 3), illumination settings (now 3) and heights (now 3), resulting in 922,057 patches extracted from 2,000 ID images (pseudo- and fully-anonymized configurations). ii) We include physical composite attacks in our novel FakeIDet2-db. To our knowledge, this is the first database that includes this kind of Presentation Attack Instrument (PAI). iii) We propose a novel approach for fake ID detection, FakeIDet2, which overcomes the limitations in patch-level fusion encountered in FakeIDet [25]. While FakeIDet performs a simple mean of scores assigned to each individual patch, FakeIDet2 comprises two learnable modules (i.e., Patch Embedding Extractor and Patch Embedding Fusion) to leverage the fusion of learned patch representations, depending on the visual characteristics from real/fake IDs. The representations at patch level are obtained through an optimized version of AdaFace [17], including dynamic class weights, which accounts for the severe class imbalance present in the databases for the different PAIs. The fusion of the representations is performed through Multi-Head Self-Attention (MHSA), which ponders the embedding values based on the relations between them, to then compress them into a single embedding, through attention pooling based on [4]. The impact of these improvements can be seen in the results (Sec. 6). And iv) In order to advance in this research field, we propose a standard public benchmark considering a wide variety of physical and synthetic attacks from different databases in the literature, allowing a direct comparison of recent approaches with the state of the art.
@@
The remainder of the paper is organized as follows. Sec. 2 provides an overview of the fake ID detection problem. In Sec. 3 we present our FakeIDet2-db, delving into the acquisition process for capturing real IDs and their fake counterparts. Sec. 4 introduces our novel privacy-aware fake ID detection method, FakeIDet2. Sec. 5 proposes a reproducible experimental protocol and standard benchmark, considering experiments in both intra- and cross-database scenarios. Finally, Sec. 6 analyzes the performance of our proposed fake detector using the proposed experimental protocol and standard benchmark and Sec. 7 draws the final conclusions.
@@ 2 Related Work
This section provides an overview of the fake ID detection problem. Sec. 2.1 provides a revision of the databases considered in the field. Sec. 2.2 focuses on representative fake ID detectors and recent international challenges. Finally, we summarize key aspects in Sec. 2.3.
@@
Table 1: Summary of databases considered in the literature for research in fake ID detection. We indicate whether the databases are available or not for public research.
@@ 2.1 Databases
Table 1 provides a summary of the most relevant public/private databases in the field of fake ID detection. One of the first family of databases was the MIDV family [1, 6, 7]. With the original purpose of Optical Character Recognition (OCR), Bulatov et al. synthetically created a set of physical fake IDs, passports, and driving licenses. They used several digital templates from multiple countries that they filled with names and addresses from Wikipedia and artificially generated faces. As different versions of the database were released, the number of fake samples increased. The DLC-2021 database [30] used the documents of the MIDV-family to create PAI s: color print, gray print, and screen. The authors also provided “real” samples, although they should be considered fake samples as they are just high-quality prints built by them, not official IDs. A similar work is presented in the KID34K dataset [28], where Park et al. introduced a dataset of 34,000 images from 82 Korean IDs and driving licenses. As in the MIDV family, the authors claimed to release “real” samples, but they were also built by the authors, not official documents. In order to prove the quality of the “real” samples, they trained a Convolutional Neural Network (CNN) on the KID34K dataset using the whole ID. The evaluation was performed using official Korean IDs and “real” documents. After that, the authors applied a dimensionality reduction technique to visualize the model’s embeddings in 2D, where they showed that embeddings from both official and “real” generated IDs from the KID34K database were similarly distributed in the embedding space. These results raise questions about the fake ID detection problem, since fake IDs can be generated with such a resemblance to real ones that deep-learning models may not be able to extract distinguishable features for proper PAD. Soares et al. [33] introduced the Brazilian Identity Document (BID) database with a total of 28,800 images from Brazilian IDs that were digitally altered using different algorithms that removed blur and replace different personal information from the real IDs to comply with data privacy regulations. In [15], Gonzalez and Tapia proposed a more sophisticated attack procedure, i.e., printing the digital copy on a Poli-Vinyl Chloride (PVC) card, which resembles even more the appearance of real IDs.
@@
Regarding synthetic data, Benalcazar et al. proposed in [5] to use Generative Adversarial Networks (GANs) to produce synthetic Chilean IDs. The proposed GAN was trained using only real Chilean ID samples. Although the data synthesized by the authors are not strictly real (and can be detected using GAN traces [26]), this was presented as a good idea to partially cover the lack of real data, for example, as a data enhancement strategy in training. Another synthetic database is IDNet [39], where Xie et al. proposed a database that comprises synthetic copies of IDs from 10 countries and 10 states of the United States of America. Given an ID template, the portrait images and the owners’ data were synthetically generated, with special care that the fields corresponded to the gender and age of the synthetic portrait image. In addition, the authors created six different types of digital manipulations, including inpainting [8], field cropping, and face morphing. A total of 837,060 IDs were generated, making it the largest and most diverse to date, although real data are not included. Markham et al. [20] developed Open-Set, a database that leverages neural style transfer, via GANs-based models to transfer textures from print and screen attacks to “real” images from the MIDV-2020 and DLC-2021 databases, creating even more samples of attacks.